duue tusdou Pagel5 SD CT and Variability For each set of data, draw and label a normal curve. each of the standard deviations above and below the mean. Specifically, label the mean and Test 1 Test 2 Test 3 154 34 54 167 132 145 17 26 34 25 14 154 145 6 24 25 23 156 154 123
Answer #1

>first test

x=c(3,7,5,4,5,6,7,8,6,5)
z=(x-mean(x))/sd(x)
> z
[1] -1.7269490 0.9298956 -0.3985267 -1.0627379 -0.3985267 0.2656845
[7] 0.9298956 1.5941068 0.2656845 -0.3985267
> d=density(z)
> plot(d)

similarly

2) ,3)

t=c(34,54,17,26,34,25,14,24,25,23)
> u=c(154,167,132,145,154,145,113,156,154,123)

z=(t-mean(t))/sd(t)
> d=density(z)
> plot(d)
> z=(u-mean(u))/sd(u)
> d=density(z)
> plot(d)
>
1)

test i.e x

density.default(x- z) CN -3 -2 0 2 N 10 Bandwidth 0.49262) test i.e tdensity.default(x- z) -2 0 N 10 Bandwidth 0.3315

3) test i.e udensity.default(x - z) CN -3 -2 0 N 10 Bandwidth 0.4748